Eliana wants to make a diorama or an art contest contest. SHE starts with a rectangular sheet of a cardboard box is 12 inches long and 8 inches wide. SHE PLANS to make the diorama box baby cute cutting identical squares from the corners and folding the sides to create a box with an open frontan open box front.
Describe how you could identify which points on a graph of v(x)=(12-2x)(8-2x)x correspond to a volume of 36 in.

Answer by stanbon(75887)

Eliana wants to make a diorama or an art contest contest. SHE starts with a rectangular sheet of a cardboard box is 12 inches long and 8 inches wide. SHE PLANS to make the diorama box baby cute cutting identical squares from the corners and folding the sides to create a box with an open frontan open box front.
Describe how you could identify which points on a graph of v(x)=(12-2x)(8-2x)x correspond to a volume of 36 in.
-----
Solve:: v(x) = 36
x(2)(6-x)(2)(4-x) = 36
x(x-6)(x-4) = 9
x(x^2-10x+24) = 9
x^3-10x^2+24x-9 = 0
------
Graph the cubic and find the roots of it.
x = 0 or x = 3 or x = 6.541..
---
Since the dimensions of the box are x, 12-2x and 8-2x
only x = 3 would keep the dimensions positive.
----
Cheers,
Stan H.
------------